OpenStudy (anonymous):

find inverse of -4cos(7x)

OpenStudy (anonymous):

to find the inverse of something I find it easiest to let y = whatever you want to invert, then rearrange for x. so here\[y=-4\cos(7x)\]\[\frac{ -y }{ 4 }=\cos(7x)\]\[\arccos(\frac{ -y }{ 4 })=7x\]\[\frac{ 1 }{ 7 }\arccos(\frac{ -y }{ 4 })=x\] So now you just replace the y with the x. So the inverse of the function you were asking is \[f^{-1}(x)= \frac{ 1 }{ 7 }\arccos(\frac{ -y }{ 4 })\]